# The Condition Number as a Scale-Invariant Proxy for Information Encoding in Neural Units

### v2

@@ 1 Introduction
The first category includes approaches such as experience replay [lopez2017gradient], which re-trains on a small portion of previous data, and knowledge distillation [li2017learning], which uses the previous model’s outputs as soft targets. Regularization techniques like Elastic Weight Consolidation (EWC) [kirkpatrick2017overcoming] and Synaptic Intelligence (SI) [zenke2017continual] also often fall into this category, as they may rely on calculating parameter importance with respect to past tasks’ data. While often effective, the practicality of these methods is significantly limited by their critical dependence on data from prior tasks, which is frequently unavailable due to privacy concerns, storage limitations, or access restrictions.
The second category comprises methods that do not require explicit access to pre-training statistics. These strategies typically focus on intrinsic model properties or architectural adjustments. Examples include architectural methods like Progressive Neural Networks (PNNs) [rusu2016progressive] and expert gateways [rosenbaum2018routing], as well as various regularization or parameter isolation techniques. Examples of the latter include Memory Aware Synapses (MAS) [aljundi2018memory] and PackNet [mallya2018packnet]. These methods attempt to preserve old knowledge by selectively modifying or freezing parts of the model without explicit access to past data.
More recently, we proposed CLASSP (Continual Learning through Adjustment Suppression and Sparsity Promotion) [ludwig2024classp], a biologically-inspired approach that mitigates forgetting by combining a decay rate over weight adjustment, i.e. suppressing updates for frequently updated weights, with a threshold on the loss gradient that promotes sparse learning by updating only weights with significant impact.
Beyond theoretical exposition, we present a practical case study where the principles of information encoding, particularly through the lens of the condition number, are applied to guide the selective fine-tuning of a multimodal Large Language Model. This application directly addresses the critical challenge of catastrophic forgetting when adapting pre-trained models to new modalities, offering a novel approach that falls into the category of methods not requiring access to prior task statistics. This demonstrates how an understanding of a tensor’s information-encoding characteristics enables more robust and efficient model adaptation strategies or even model pruning [ludwig2023compressing].
@@ 2 From Stability to Semantics: Rethinking the Condition Number
Beyond their well-established role in numerical stability and optimization, the singular values and condition number of neural network weight matrices have also become a focal point for understanding model behavior, generalization, and information processing. Prior work has shown that a flatter singular value spectrum can promote better generalization by enabling more uniform transformations and reducing sensitivity to input perturbations [jia2017improving, ludwig2014eigenvalue]. Other studies have explored how singular values relate to information compression and flow through network layers, aligning with the principles of the Information Bottleneck framework [tishby2015deep].
@@ 3 Information Encoding and Entropy
3.1 Measures of Information Encoding: Scale Invariance
The concept of mutual information $I(X;Y)$ quantifies the amount of information that $X$ and $Y$ share, or how much knowing one of these variables reduces uncertainty about the other. For continuous variables, it is defined as:
I(X;Y)=h(X)h(X|Y)=h(Y)h(Y|X)I(X;Y)=h(X)-h(X|Y)=h(Y)-h(Y|X) (7)
For a deterministic transformation $Y=WX$, if $X$ is known, $Y$ is also known, meaning $h(Y|X)=0$. In this case, the mutual information simplifies to $I(X;Y)=h(Y)$.
Note that that differential entropy $h(Z)$ is a scale-variant measure. If a continuous random variable $Z$ is scaled by a constant $c$, its differential entropy changes by $h(cZ)=h(Z)+\log_{2}|c|$. Consequently, for a deterministic linear transformation, $I(X;Y)=h(Y)$ also inherits this scale-variance. Therefore, interpreting the absolute values of $h(Y)$ or $I(X;Y)$, in this deterministic continuous context, as a direct, unit-independent measure of information content or uncertainty reduction requires careful consideration of the absolute scaling factors.
To characterize how the weight tensor $W$ itself processes and encodes information, we initially considered two candidate measures derived from its singular values: the condition number and the log-volume scaling factor. The condition number $\kappa(W)$ (1) reflects the anisotropy, i.e. relative stretching/compression, of the transformation. The condition number is scale-invariant with respect to $W$; if $W$ is scaled by a non-zero scalar $c$, $\kappa(cW)=\kappa(W)$, and is not a function of $X$ (no data required). The log-volume scaling factor of the transformation, last term of (6), represents the logarithm of the absolute factor by which the transformation $W$ scales volume in the relevant subspace. Although this measure is directly related to entropy, it is sensitive to the overall scale of $W$.
Given our objective to analyze intrinsic properties of the weight tensor’s transformation independent of arbitrary scaling, the condition number $\kappa(W)$ was chosen as our primary measure due to its property of being invariant to the scale of $W$. This allows us to focus purely on the relative distribution of singular values and the anisotropic nature of the learned transformation, which is key to understanding selective information encoding. The log-volume scaling factor, while informative about overall volume change, is not suitable for a scale-invariant analysis of the weight tensor’s inherent encoding strategy.
@@ 3.2 A Proxy for Knowledge Encoding: Scale-Invariant Perspective
As proven, a uniform distribution of singular values maximizes the log-volume scaling factor for a given Frobenius norm, leading to increased information transfer with higher output entropy and more uncertainty about the output vector. Conversely, tensors with non-uniform singular value distributions (higher condition numbers) exhibit anisotropic behavior, selectively amplifying certain directions while compressing others. In well-trained models, this anisotropy reflects learned feature hierarchies that concentrate representational capacity on task-relevant signals while attenuating less predictive input variations, reducing output entropy while preserving predictive power. This perspective aligns with the Information Bottleneck principle [tishby2015deep], where effective representations compress input information, reducing mutual information $I(X;Z)$ while retaining task-relevant features, maximizing mutual information $I(Z;Y)$, where $X$ denotes the input data, $Z$ the compressed learned representation, and $Y$ the target output.
@@ 4.1 KappaTune: General Algorithm
4.2 Adapting a pre-trained LLM to a new task (MoE sarcasm detection)
KappaTune is further assessed on OLMoE [muennighoff2024olmoe], an open-source MoE-based LLM (1B active parameters, 7B total), configured with 64 experts per layer and top-8 expert activation. The motivation for using an MoE model is the finer granularity for tensor selection in modular experts [^3] https://github.com/oswaldoludwig/kappaTune/blob/main/experiments_sarcasm_kappatune.py.
@@ 4.3 Adapting a pre-trained LLM to a new input modality (ASR)
The MM-LLM is composed with a frozen audio encoder and a text LLM connected by a trainable adaptor. The audio encoder is the Whisper large-v2 encoder (frozen) [radford2022robustspeechrecognitionlargescale], which transforms raw audio into high-dimensional embeddings. The text LLM is Llama-3.2-8B-Instruct [llama3-herd]. The trainable adaptor bridges Whisper outputs to the Llama embedding space via a linear projection, layer normalization, and two multihead attention layers with positional encoding. Audio features are sub-sampled $4\times$ by concatenation.

### v3

@@ 1 Introduction
The first category includes approaches such as experience replay [8], which re-trains on a small portion of previous data, and knowledge distillation [6], which uses the previous model’s outputs as soft targets. Regularization techniques like Elastic Weight Consolidation (EWC) [5] and Synaptic Intelligence (SI) [19] also often fall into this category, as they may rely on calculating parameter importance with respect to past tasks’ data. While often effective, the practicality of these methods is significantly limited by their critical dependence on data from prior tasks, which is frequently unavailable due to privacy concerns, storage limitations, or access restrictions.
The second category comprises methods that do not require explicit access to pre-training statistics. These strategies typically focus on intrinsic model properties or architectural adjustments. Examples include architectural methods like Progressive Neural Networks (PNNs) [17] and expert gateways [16], as well as various regularization or parameter isolation techniques. Examples of the latter include Memory Aware Synapses (MAS) [2] and PackNet [13]. These methods attempt to preserve old knowledge by selectively modifying or freezing parts of the model without explicit access to past data.
More recently, we proposed CLASSP (Continual Learning through Adjustment Suppression and Sparsity Promotion) [11], a biologically-inspired approach that mitigates forgetting by combining a decay rate over weight adjustment, i.e. suppressing updates for frequently updated weights, with a threshold on the loss gradient that promotes sparse learning by updating only weights with significant impact.
Beyond theoretical exposition, we present a practical case study where the principles of information encoding, particularly through the lens of the condition number, are applied to guide the selective fine-tuning of a multimodal Large Language Model. This application directly addresses the critical challenge of catastrophic forgetting when adapting pre-trained models to new modalities, offering a novel approach that falls into the category of methods not requiring access to prior task statistics. This demonstrates how an understanding of a tensor’s information-encoding characteristics enables more robust and efficient model adaptation strategies or even model pruning [9].
@@ 2 From Stability to Semantics: Rethinking the Condition Number
Beyond their well-established role in numerical stability and optimization, the singular values and condition number of neural network weight matrices have also become a focal point for understanding model behavior, generalization, and information processing. Prior work has shown that a flatter singular value spectrum can promote better generalization by enabling more uniform transformations and reducing sensitivity to input perturbations [4, 10]. Other studies have explored how singular values relate to information compression and flow through network layers, aligning with the principles of the Information Bottleneck framework [18].
@@ 3 Information Encoding and Entropy
3.1 Geometric Foundations and Scale-Invariance
Consider the aforementioned linear transformation $Y=WX$, where $W\in\mathbb{R}^{m\times n}$. The singular values $\sigma_{i}$ of $W$ govern the axis-wise scaling of the transformation. While the Frobenius norm $\|W\|_{F}$ reflects the overall scaling of the transformation, the condition number $\kappa(W)=\sigma_{\max}/\sigma_{\min}$ (for full-rank $W$) characterizes the anisotropy of the transformation and is invariant to uniform scaling of $W$.
This scale-invariant geometric property represents the unit’s directional selectivity. A unit with $\kappa(W)\approx 1$ preserves the geometry of the input space, whereas a unit with high $\kappa(W)$ prioritizes a specific subspace while suppressing others. As shown in the next subsections, this anisotropy becomes a key determinant of the unit’s information encoding behavior under fixed-norm constraints.
@@ 3.2 A Proxy for Knowledge Encoding: Scale-Invariant Perspective
As proven, a uniform distribution of singular values maximizes the log-volume scaling factor for a given Frobenius norm [^1] The differential entropy $h(Z)$ is a scale-variant measure. For a random vector $Z\in\mathbb{R}^{m}$ and constant $c\neq 0$, $h(cZ)=h(Z)+\log_{2}|c|$, leading to increased information transfer with higher output entropy and more uncertainty about the output vector. Conversely, tensors with non-uniform singular value distributions (higher condition numbers) exhibit anisotropic behavior, selectively amplifying certain directions while compressing others. In well-trained models, this anisotropy reflects learned feature hierarchies that concentrate representational capacity on task-relevant signals while attenuating less predictive input variations, reducing output entropy while preserving predictive power. This perspective aligns with the Information Bottleneck principle [18], where effective representations compress input information, reducing mutual information $I(X;Z)$ while retaining task-relevant features, maximizing mutual information $I(Z;Y)$, where $X$ denotes the input data, $Z$ the compressed learned representation, and $Y$ the target output.
@@
In summary, Theorem 1 establishes, under the given assumptions, that tensors with low condition numbers maximize differential entropy, positioning them as the most information-rich and potentially least specialized components. This theoretical foundation explains their suitability for adaptation tasks, where broad representational capacity facilitates efficient learning without disrupting pre-existing information encoding.
@@ 4.1 KappaTune: General Algorithm
4.2 Adapting a pre-trained LLM to a new task
KappaTune is assessed on OLMoE [14], an open-source MoE-based LLM (1B active parameters, 7B total), configured with 64 experts per layer and top-8 expert activation. The motivation for using an MoE model is the finer granularity for tensor selection in modular experts [^4] https://github.com/oswaldoludwig/kappaTune/blob/main/experiments_sarcasm_kappatune.py.
@@ 4.2 Adapting a pre-trained LLM to a new task
4.3 Adapting a pre-trained LLM to a new task via KappaTune-LoRA
LoRA adapters enable task-specific fine-tuning while allowing attachment and detachment to avoid catastrophic forgetting. However, preserving pre-trained general knowledge, even when LoRA adapters are attached for a specific task, enhances the model’s reasoning capabilities on the target task, thereby improving generalization capacity. To assess this in a high-parameter regime, we perform a comparative analysis on the DeepSeek-V2-Lite LLM [7], a MoE model with 16B total parameters and $\sim$2.4B active parameters.
Specifically, we evaluate whether KappaTune-LoRA, which strategically places LoRA adapters on tensors with low condition numbers, outperforms standard uniform placement. This comparison controls the trainable parameters by approximately 190 million for both experiments, focusing on generalization in sentiment analysis [^6] https://github.com/oswaldoludwig/kappaTune/blob/main/experiments_SA_kappatune.py.
The model is loaded in 4-bit precision and two adaptation strategies are adopted:
  • LoRA Global: Standard low-rank adaptation ($r=16$, $\alpha=32$) targeting all major linear projections (q_proj, v_proj, up_proj, down_proj). This configuration serves as the dense adaptation baseline.
  • KappaTune-LoRA: The condition number ($\kappa$) is computed for all expert tensors in the MoE layers. A budget of $K=300$ tensors with the lowest $\kappa$ is selected. To ensure a fair comparison of capacity, the rank is increased to $r=190$ for these selected adapters, resulting in the same number of trainable parameters as LoRA Global.
[…]
Table 2: Generalization and Forgetting Analysis. Lower values are better for both metrics.
[…]
($\text{PPL}_{\text{IMDB Test}}-\text{PPL}_{\text{IMDB Train}}$) ($\text{PPL}_{\text{Wiki Fine-Tuned model}}-\text{PPL}_{\text{Wiki Base model}}$)
LoRA Global +2.06 +3.44
KappaTune +1.61 +1.21
[…]
This superior generalization supports our hypothesis that by restricting adaptation to stable and less specialized tensors, KappaTune preserves the model’s broad linguistic capabilities (forgetting cost of only +1.21). This retained general knowledge acts as an inductive bias, aiding the model in reasoning about the new task rather than relying solely on surface-level pattern matching from the training data.
@@ 4.4 Adapting a pre-trained LLM to a new input modality (ASR)
The MM-LLM is composed with a frozen audio encoder and a text LLM connected by a trainable adaptor. The audio encoder is the Whisper large-v2 encoder (frozen) [15], which transforms raw audio into high-dimensional embeddings. The text LLM is Llama-3.2-8B-Instruct [1]. The trainable adaptor bridges Whisper outputs to the Llama embedding space via a linear projection, layer normalization, and two multihead attention layers with positional encoding. Audio features are sub-sampled $4\times$ by concatenation.